# Vacancy-Induced Quantum Properties in 2D Silicon Carbide: Atomistic insights from semi-local and hybrid DFT calculations

### v1

@@ IV Conclusions
This comprehensive study of silicon and carbon vacancies in monolayer SiC (1L-SiC) using density functional theory (DFT) has provided crucial insights into the electronic, structural, and optical properties of these defects. Our investigation, employing a range of exchange-correlation functionals (PBE, SCAN, r2SCAN, and HSE06) and charge correction schemes (FNV and KO), demonstrates the critical importance of methodology in accurately predicting defect properties in 2D materials. The formation energies of silicon and carbon vacancies reveal significant variations across functionals, with HSE06 consistently predicting the highest values, highlighting the substantial impact of exact exchange inclusion on defect energetics. The electronic structure analysis demonstrates distinct characteristics between silicon and carbon vacancies. Silicon vacancies introduce highly localized defect states, with HSE06 calculations revealing strong spin polarization effects and a gap of approximately 2 eV for spin-up states. In contrast, carbon vacancies exhibit more dispersed states with broader bandwidth ( 0.5 eV), suggesting stronger orbital overlap with neighboring atoms. The progression of electronic structure across charge states shows intriguing magnetic behavior, particularly evident in HSE06 calculations, where spin polarization persists even in charged states. Migration pathway analysis reveals significantly lower barriers for silicon vacancies (0.38 eV) compared to carbon vacancies (0.81 eV), indicating higher mobility of silicon vacancies at moderate temperatures. This finding has important implications for defect dynamics and the potential formation of vacancy complexes in 1L-SiC. Our optical property calculations using PBE-DFPT reveal significant modifications to the dielectric response and absorption characteristics, particularly in the far-infrared region (0-0.5 eV). Positively charged vacancies demonstrate the strongest optical response, with absorption coefficients reaching 22%. The distinct behavior of silicon and carbon vacancies - with silicon vacancies showing sharp, discrete absorption features and carbon vacancies exhibiting broader absorption profiles - provides complementary functionalities for potential optoelectronic applications. The choice of computational methodology proves crucial for accurate predictions. While HSE06 provides the most reliable results for electronic structure and energetics, the computational efficiency of meta-GGA functionals like SCAN and r2SCAN makes them attractive for initial screening studies. The PBE-DFPT approach demonstrates reliability for optical property calculations while maintaining computational efficiency. These findings open several promising directions for future research. The strong charge-state dependence of optical properties suggests possibilities for electrostatically tunable optical devices operating in the far-infrared region. The distinct characteristics of silicon and carbon vacancies could be exploited for dual-functionality devices combining sharp, state-specific optical transitions with broader absorption bands. Future work should focus on exploring the interplay between different types of defects, investigating strain effects on defect properties, and experimental validation of these theoretical predictions. This work establishes a comprehensive framework for investigating defects in 2D materials and demonstrates how careful selection of computational methods can provide detailed insights into defect properties. The complementary functionalities of silicon and carbon vacancies revealed in this study suggest exciting possibilities for defect engineering in next-generation quantum and optoelectronic devices based on 2D SiC.
V Author contributions
@@
VI Conflicts of interest
There are no conflicts to declare.

### v2

@@ 4 Conclusions
This comprehensive study of silicon and carbon vacancies in monolayer SiC (1L-SiC) using density functional theory has provided crucial insights into the electronic, structural, and optical properties of these defects. Our investigation, employing a range of exchange-correlation functionals (PBE, SCAN, r2SCAN, and HSE06) and charge correction schemes (FNV and KO), demonstrates the critical importance of methodology in accurately predicting defect properties in 2D materials. The formation energies of silicon and carbon vacancies reveal significant variations across functionals, with HSE06 consistently predicting the highest values (7.5-8.5 eV for Si vacancies and 6.0-7.0 eV for C vacancies at the VBM), highlighting the substantial impact of exact exchange inclusion on defect energetics. The relative stability of carbon vacancies over silicon vacancies by 1.0-1.5 eV persists across all functionals, providing a robust prediction independent of methodology. The charge transition levels, while differing quantitatively across functionals, maintain consistent relative positions, with the +1/0 and 0/-1 transitions occurring at approximately 1.0 eV and 2.2 eV above the VBM, respectively, for both vacancy types. The electronic structure analysis demonstrates distinct characteristics between silicon and carbon vacancies. Silicon vacancies introduce highly localized defect states, with HSE06 calculations revealing strong spin polarization effects and a gap of approximately 2 eV for spin-up states. In contrast, carbon vacancies exhibit more dispersed states with broader bandwidth ( 0.5 eV), suggesting stronger orbital overlap with neighboring atoms. This fundamental difference in localization behavior has important implications for the potential applications of these defects in quantum technologies. The highly localized nature of silicon vacancy states makes them promising candidates for quantum bit (qubit) applications, potentially offering longer coherence times due to their reduced interaction with the environment, while carbon vacancies might be more suitable for applications requiring enhanced coupling to external fields. The progression of electronic structure with changing charge states follows systematic trends, with HSE06 calculations revealing significant spin polarization across all charge states for both vacancy types. The largest spin splitting of 1.3 eV is observed for V, suggesting that this configuration could exhibit the strongest magnetic properties among the studied systems. The orbital character analysis reveals predominantly p-orbital contributions for silicon vacancy states, while carbon vacancies show mixed s-p character, further explaining their different electronic properties. Our migration pathway analysis reveals significantly lower barriers for silicon vacancies (0.8 eV for the most favorable path) compared to carbon vacancies (1.2 eV), indicating higher mobility of silicon vacancies at moderate temperatures. This finding has important implications for defect dynamics and the potential formation of vacancy complexes in 1L-SiC, particularly under thermal annealing conditions where Si vacancy diffusion would dominate. The detailed transition state analysis provides atomic-level insights into the migration mechanisms, revealing that the stronger covalent bonding character of carbon atoms leads to more pronounced structural distortions during migration, explaining their higher barriers. The optical property calculations reveal significant modifications to the dielectric response and absorption characteristics, particularly in the far-infrared region (0-0.5 eV). Positively charged vacancies demonstrate the strongest optical response, with absorption coefficients reaching 22% enhancement relative to pristine SiC. The distinct behavior of silicon vacancies (showing sharp, discrete absorption features) and carbon vacancies (exhibiting broader absorption profiles) provides complementary functionalities for potential optoelectronic applications. These optical signatures could serve as spectroscopic fingerprints for identifying and characterizing defects in experimental samples. The choice of computational methodology proves crucial for accurate predictions. While HSE06 provides the most reliable results for electronic structure and energetics among the functionals used, even more accurate methods, such as GW or diffusion Monte Carlo, exist but are computationally prohibitive for the periodic systems and large-scale defect calculations in this study 46. For molecular systems, coupled cluster methods like CCSD(T) serve as a gold standard, but their application to periodic crystals like 2D SiC is impractical. The computational efficiency of meta-GGA functionals like SCAN and SCAN makes them attractive for initial screening studies. The PBE-DFPT approach demonstrates reliability for optical property calculations while maintaining computational efficiency. The importance of appropriate charge correction schemes is underscored by the significant differences (up to 0.5 eV) between the FNV and KO methods, with the latter’s treatment of anisotropic dielectric screening being particularly important for 2D materials. These findings open several promising directions for future research. The strong charge-state dependence of optical properties suggests possibilities for electrostatically tunable optical devices operating in the far-infrared region. The distinct characteristics of silicon and carbon vacancies could be exploited for dual-functionality devices combining sharp, state-specific optical transitions with broader absorption bands. The significant difference in mobility between Si and C vacancies suggests potential routes for controlled formation of vacancy complexes through selective thermal annealing. Future work should focus on exploring the interplay between different types of defects, investigating strain effects on defect properties, and experimental validation of these theoretical predictions. Studies of defect complexes, particularly the divacancy (V-V) and transition metal dopant-vacancy pairs, could reveal novel quantum functionalities. Additionally, exploring the impact of external fields (electric, magnetic, strain) on defect properties could provide routes for dynamic control of quantum states. This work establishes a comprehensive framework for investigating defects in 2D materials and demonstrates how careful selection of computational methods can provide detailed insights into defect properties. The complementary functionalities of silicon and carbon vacancies revealed in this study suggest exciting possibilities for defect engineering in next-generation quantum and optoelectronic devices based on 2D SiC.
The author acknowledges the computational resources provided by the University of Delaware High Performance Computing facility. Special thanks to colleagues at the Delaware Energy Institute for valuable discussions.
Supplementary information containing convergence test results anddetailed analysis of density of states (DOS) for all vacancy configurations, additional band structure plots, and transition state geometries for vacancy migration is available free of charge via the Internet at http://pubs.acs.org.
5 Author Contributions
@@
6 Conflicts of Interest
The author declares no competing financial interest.